A committee for the end-of-year party is composed of four eighth graders and three seventh graders. A three-member subcommittee is formed...
Find the probability that all 3 members on the subcommittee are eighth graders?
Thanks for your help

Answer by stanbon(75887)

A committee for the end-of-year party is composed of four eighth graders and three seventh graders. A three-member subcommittee is formed...
Find the probability that all 3 members on the subcommittee are eighth graders?
-------------------
# of ways to pick 3 8th graders: 4C3 = 4
# of ways to pick 3 anything: 7C3 = 35
-------------------------------------
P(3 are 8th) = 4C3/7C3 = 4/35 = 0.1143
=============
Cheers
Stan H.